# On the $\ell$-th largest degree of an intersecting family

### v1

@@ 4 Estimating the $(\ell+1)$-th largest degree
We split the set in $\mathcal{F}$ that contains $i$ into three parts and estimate the size of each part separately.
$D_{1}=\{A|i\in A\in\mathcal{F},|A\cap[\ell+1]|>m\},$
$D_{2}=\{A|i\in A\in\mathcal{F},|A\cap[\ell+1]|\leq m,A\cap[\ell+1]\not\in\mathcal{G}\},$
$D_{3}=\{A|i\in A\in\mathcal{F},A\cap[\ell+1]\in\mathcal{G}\}.$
(a) For $D_{1}$, consider those $A$ satisfy $i\in A\in\mathcal{F}$ and $|A\cap[\ell+1]|\geq m+1$. Similar to the proof of Theorem 1.4, we bound the number of $|D_{1}|$ by:
$|D_{1}|\leq\sum_{j=m+1}^{\ell+1}\binom{\ell}{j-1}\binom{n-\ell-1}{k-j}\leq 2^{\ell}\binom{n-\ell-1}{k-m-1}$.
(b) For $D_{2}$, those $A$ satisfy $i\in A\in\mathcal{F}$, $|A\cap[\ell+1]|\leq m$, but $A\cap[\ell+1]\notin\mathcal{G}$. The number of such $A$ can be bounded by:
$|D_{2}|\leq 2^{\ell}\binom{n-\ell-1}{k-\ell-2+m}$.
(c) For $D_{3}$, consider those $A$ with $i\in A\in\mathcal{F}$ with $A\cap[\ell+1]\in\mathcal{G}$. Then:
$|D_{3}|$ $\leq d_{\{1,i\}}(\mathcal{F})+\sum_{j=3}^{m}\binom{\ell-1}{j-2}\binom{n-\ell-1}{k-j}$
$\leq\binom{l-1}{0}\binom{n-l-1}{k-2}-\binom{n-k-2}{k-2}+\sum_{j=3}^{k}\binom{\ell-1}{j-2}\binom{n-\ell-1}{k-j}$
$=\binom{n-2}{k-2}-\binom{n-k-2}{k-2}$.
Now we show that, under the conditions $n>\dfrac{k^{2}}{\ell}+64k$ and $k\geq\ell\geq 20$,
$\binom{n-k-2}{k-2}\geq 2^{\ell+1}\binom{n-\ell-1}{k-m-1}>|D_{1}|+|D_{2}|.$
$\frac{\binom{n-k-2}{k-2}}{\binom{n-\ell-1}{k-m-1}}$ $=\prod_{j=1}^{k-m-1}\frac{n-k-1-j}{n-\ell-j}\;\prod_{j=1}^{m-1}\frac{n-2k+j}{k-m-1+j}$
$\geq\left(1-\frac{\ell}{k+62\ell}\right)^{k-m-1}\left(\frac{k}{\ell}+62\right)^{m-1}$
[…]
$\geq 2^{\ell+1}.$
The first inequality is because
$\dfrac{n-k-1-j}{n-\ell-j}=1-\dfrac{k+1-\ell}{n-\ell-j}\geq 1-\dfrac{k}{\frac{k^{2}}{\ell}+64k-\ell-k}\geq 1-\frac{\ell}{k+62\ell},$
$\dfrac{n-2k+j}{k-m-1+j}\geq\dfrac{\frac{k^{2}}{\ell}+64k-2k+m-1}{k-2}\geq\dfrac{k}{\ell}+62$
Therefore,
$d_{i}(\mathcal{F})\leq|D_{1}|+|D_{2}|+|D_{3}|\leq\binom{n-2}{k-2}<\binom{n-2}{k-2}+\binom{n-\ell-1}{k-\ell}$.
@@
Case 2: $d_{i}^{(2)}(\mathcal{G})=1$ and $d_{\{1,i\}}(\mathcal{F})>\binom{n-\ell-1}{k-2}-\binom{n-k-2}{k-2}$. Then, by the same argument as in the proof of Theorem 1.4, we have $d_{j}^{(u)}(\overline{\mathcal{G}})\leq\binom{\ell-1}{u-2}$ for all $j\neq 1$ and $u\leq m$. Consequently, if for some $j\neq 1$ we have $d_{\{1,j\}}(\mathcal{F})\leq\binom{n-\ell-1}{k-2}-\binom{n-k-2}{k-2}$, then the proof is already complete (by reducing to the situation of Case 1). Thus we may assume that for every $j\neq 1$,
@@ 5 Concluding Remarks
In this paper, we study the problem of determining the maximum value of the $(\ell+1)$-th largest degree of an intersecting family $\mathcal{F}\subset\binom{[n]}{k}$. For $\ell=2k$, Theorem 1.2 provides the precise range that the $1$-star gives the maximum, resolving the Frankl–Wang Conjecture. For $\ell=k+1$, Theorem 1.4 shows that the $1$-star example gives the maximum for $n>\frac{11}{2}k$. A natural question is whether the constant $\frac{11}{2}$ can be reduced to $2$.

### v3

@@ 4 Estimating the $(\ell+1)$-th largest degree
We split the set in $\mathcal{F}$ that contain $i$ into the following two parts and estimate the size of each part separately:
$D_{1}=\{A|i\in A\in\mathcal{F},A\cap[\ell+1]\in\mathcal{G}\}$
$D_{2}=\{A|i\in A\in\mathcal{F},A\cap[\ell+1]\not\in\mathcal{G}\}$
$|D_{1}|$ can be bounded using the inequalities $d_{i}^{(j)}(\overline{\mathcal{G}})\leq\binom{\ell-1}{j-2}$:
$|D_{1}|$ $=\sum_{i\in X\in\mathcal{G}}d_{X}(\mathcal{F})$
$\leq\sum_{j\geq 3}d_{i}^{(j)}(\mathcal{G})\binom{n-\ell-1}{k-j}+d_{\{1,i\}}(\mathcal{F})$
$\leq\sum_{j\geq 3}d_{i}^{(j)}(\overline{\mathcal{G}})\binom{n-\ell-1}{k-j}+d_{\{1,i\}}(\mathcal{F})$
$\leq\sum_{j\geq 3}\binom{\ell-1}{j-2}\binom{n-\ell-1}{k-j}+\binom{n-\ell-1}{k-2}-\binom{n-k-2}{k-2}$
$=\binom{n-2}{k-2}-\binom{n-k-2}{k-2}$
Next we estimate $|D_{2}|$. We claim that for $X\not\in\mathcal{G}$, we have $d_{X}(\mathcal{F})\leq 2^{k}\binom{n-k}{k-m}$.
If $|X|\leq m$, then the inequality holds follows from the definition of $\mathcal{G}$. If $|X|>m$, then
$d_{X}(\mathcal{F})$ $\leq\binom{n-\ell-1}{k-m}\leq\sum_{j=0}^{k-m}\binom{n-k}{j}\binom{k-\ell-1}{k-m-j}\leq 2^{k-\ell-1}\binom{n-k}{k-m}\leq 2^{k}\binom{n-k}{k-m}$
Hence $|D_{2}|$ can be bounded by the following inequality,
$|D_{2}|$ $\leq\sum_{i\in X\not\in\mathcal{G}}d_{X}(\mathcal{F})\leq 2^{\ell+1}\cdot 2^{k}\binom{n-k}{k-m}$.
We want to show that, there exists some constant $C_{\varepsilon}$, such that when $\ell\geq\varepsilon k$ and $n\geq C_{\varepsilon}k$,
$2^{\ell+k+1}\binom{n-k}{k-m}\leq\binom{n-k-2}{k-2}$.
In fact, we can derive the following:
$\dfrac{\binom{n-k-2}{k-2}}{\binom{n-k}{k-m}}$ $\geq\left(\dfrac{n-2k}{k}\right)^{m-2}\cdot\left(\dfrac{n-2k}{n-k}\right)^{2}\geq 2^{-2}\left(\dfrac{n}{k}-2\right)^{\frac{\ell}{2}-3}.$
When $C_{\varepsilon}\geq 3^{2+\frac{2}{\varepsilon}}>3^{2+\frac{2k}{\ell}}$, for sufficiently large $k$, when $n\geq C_{\varepsilon}k$,
$\dfrac{\binom{n-k-2}{k-2}}{\binom{n-k}{k-m}}\geq 2^{-2}\left(\dfrac{n}{k}-2\right)^{\frac{\ell}{2}-3}\geq 2^{-2}\left(3^{2+\frac{2k}{\ell}}\right)^{\frac{\ell}{2}-3}\geq 2^{k+\ell+1}.$
Where the last inequality holds when $k$ is sufficiently large.
Therefore, taking $C_{\varepsilon}=3^{2+\frac{2}{\varepsilon}}$, in Case 1 we have
$d_{i}(\mathcal{F})\leq\binom{n-2}{k-2}\leq\binom{n-2}{k-2}+\binom{n-\ell-1}{k-\ell}$,
which is what we wanted.
@@
Case 2: $d_{i}^{(2)}(\overline{\mathcal{G}})=1$ and $d_{\{1,i\}}(\mathcal{F})>\binom{n-\ell-1}{k-2}-\binom{n-k-2}{k-2}$. Then, by the same argument as in the proof of Theorem 1.4, we have $d_{j}^{(u)}(\overline{\mathcal{G}})\leq\binom{\ell-1}{u-2}$ for all $j\neq 1$ and $u\leq m$. Consequently, if for some $j\neq 1$ we have $d_{\{1,j\}}(\mathcal{F})\leq\binom{n-\ell-1}{k-2}-\binom{n-k-2}{k-2}$, then the proof is already complete (by reducing to the situation of Case 1). Thus we may assume that for every $j\neq 1$,
@@ 5 Concluding Remarks
In this paper, we study the problem of determining the maximum value of the $(\ell+1)$-th largest degree of an intersecting family $\mathcal{F}\subset\binom{[n]}{k}$. For $\ell=2k$, Theorem 1.2 provides the precise range that the $1$-star gives the maximum, resolving the Frankl–Wang Conjecture. For $\ell=k+1$, Theorem 1.4 shows that the $1$-star example gives the maximum for $n>12k$. A natural question is whether the constant $12$ can be reduced.
@@
Acknowledgment. The authors thank Yongtao Li for pointing out a misstated theorem used in the previous version of this paper.